A rectangular swimming pool that is 7 meters deep, 11 meters long, and 10 meters wide. water is $20 per kiloliter. How much will it cost to fill the pool?
Here is what I have written, just don't remember how to set it up properly.
7*11*10=volume*1 kl/m^3*20 $/kl=$15,400

Answer by KMST(5396)

In steps, it could be:
Volume = (7m)(11m)(10m) = 770m^3
Volume = (770m^3)(1 kL/m^3) = 770 kL, for the conversion to kiloliters
Price = (770 kL)($20/kL) = $15,400
All in one line:
Price = (7m)(11m)(10m)(1 kL/m^3)($20/kL) = $15,400